OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

x = 5/3, x = - 10/3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (dmndlife24):

-3x -21 = -12 -3x = 9 x = -3

OpenStudy (texaschic101):

-3 | x + 7 | = -12 --- we want to get the absolute value and everything in it on one side by itself...so divide both sides by -3 | x + 7 | = - 4 x + 7 = -4 x = -4 - 7 x = - 11 -(x + 7) = -4 -x - 7 = -4 -x = -4 + 7 -x = 3 x = -3 so x = -11 and x = -3

OpenStudy (anonymous):

Great, thank you!! Do you think you could help me with some other questions?

OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 |5x − 6| = −41 | Equation 2 |7x + 13| = 27

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

The first drawing is equation 1, second is equation 2

OpenStudy (anonymous):

Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

I hope someone can help!

OpenStudy (texaschic101):

| 5x - 6 | = -41 5x - 6 = -41 -(5x - 6) = -41 5x = -41 + 6 -5x + 6 = -41 5x = 35 -5x = -41 - 6 x = 7 -5x = - 47 x = 47/5 =========================== | 7x + 13| = 27 7x + 13 = 27 -(7x + 13) = 27 7x = 27 - 13 -7x - 13 = 27 7x = 14 -7x = 27 + 13 x = 2 -7x = 40 x = - 40/7 ============================= well....they each have 2 solutions...so I am going with second answer choice